# Generalization capabilities and robustness of hybrid machine learning models grounded in flow physics compared to purely deep learning models

### v2

@@ II.2 Probabilistic forecasting: Variational autoencoder
In contrast to point forecasting, our goal here is to approximate the probability distribution $\mathbb{F}_{\bm{v}}$ that governs the forecasting process from a sequence of previous snapshots $\{\bm{v}_{i-n+1},\bm{v}_{i-n+2},\dots,\bm{v}_{i}\}$ to the next snapshot $\{\bm{v}_{i+1}\}$. To achieve this, we define a third model which adopts the Bayesian perspective of statistical inference and develop a variational inference model (Blei, Kucukelbir, and McAuliffe, 2017) based on variational autoencoders (VAEs) (Kingma and Welling, 2022). This model is presented in this article for the first time, to the authors’ knowledge. For a visual depiction of the proposed model, see Figure 4.
Figure 4: The variational autoencoder proposed in this work. Note that here both the exact and approximate posteriors, $p_{0}(\bm{z}|\cdot)$ and $q(\bm{z}|\cdot)$, respectively, are conditioned on the input sequence of snapshots while we compute the likelihood $p(\bm{v}_{t+1}|\bm{z})$ of the time-ahead snapshot conditioned on the latent variables.
As with the residual autoencoder, VAE adopts an encoder-decoder structure that operate on the spatial dimension of the snapshots, and there is a hidden block that operates on the temporal dimension to capture the temporal correlations within the latent space. The specific architecture used for the laminar flow is detailed in Tables 9 and 10.
@@
Note from Figure 4 that the decoder input comes from the posterior distribution $p_{0}(\bm{z}|\{\bm{v}_{i}\}_{i=t-n+1}^{t})$. Therefore, it is necessary to generate a sample from this posterior and use it as input to the decoder. The same is done with the likelihood $p(\bm{v}_{t+1}|\bm{z})$ to obtain the time-ahead predicted snapshot.
@@
Where $\bm{x}$ is the input image, $p_{0}(\bm{z}|\bm{x})$ is the exact posterior of the latent variables $\bm{z}$, $q(\bm{z}|\bm{x})$ is the approximate posterior, and $p(\bm{x}|\bm{z})$ is the likelihood of the image $\bm{x}$ given the latent variables $\bm{z}$. Note that in the original formulation (4), the ELBO is evaluated for a single image $\bm{x}$.
In this work, we explore a modification of this loss function, where the exact and approximate posterior are conditioned on the input sequence of snapshots $\{\bm{v}_{i}\}_{i=t-n+1}^{t}$ instead of a single image $\bm{x}$. Additionally, we aim to approximate the likelihood of the time-ahead snapshot $p(\bm{v}_{t+1}|\bm{z})$. Therefore the loss function used to train the VAE is as follows,
@@
This loss function allows to approximate the likelihood of the time-ahead snapshot based on the latent variables $\bm{z}$ that are obtained from the previous sequence of snapshots $\{\bm{v}_{i}\}_{i=t-n+1}^{t}$.
@@ III Results
After extensive experimentation, it was observed that the residual autoencoder and VAE models perform better when tasked with predicting a single time-ahead sample, in contrast to the POD-DL model. The architectural simplicity of the POD-DL model enables it to leverage a smaller training set while still predicting multiple time-ahead predictions effectively, i.e., a longer horizon. The number of training and testing samples for each dataset is shown in Table 3.
@@
Figure 5: Comparison of the ground truth streamwise velocity evolution $u_{x}(t)$, at spatial coordinate $P_{1}$ shown in Figure 9, against the predictions obtained from the (a) POD-DL, (b) residual autoencoder and (c) variational autoencoder models, respectively. Counter part for the spatial coordinate $P_{2}$ in figures (d), (e) and (f). In all figures the ground truth is represented by a solid line and the prediction by a dashed line.
@@ III.1 Three-dimensional wake behind a circular cylinder: laminar flow
This dataset represents a three-dimensional wake of a circular cylinder, from a viscous, incompressible, and Newtonian flow, which is described by the Navier–Stokes equations, written in non-linear form, as follows,
@@
The first $100$ snapshots of the simulation correspond to the transitory state of the flow, as is shown in Figure 6. It is known that in this regime the flow statistics are constantly varying, thus these first snapshots are not considered for training and testing the models. Figure 6 shows how snapshots from $100$ to $399$ are used to train the models, while the rest ($400$ to $599$) are used to testing.
@@
The decay of the $50$ most energetic modes and cummulative energy $E$ are shown in Figure 8. The cummulative energy is defined as follows,
@@
where $\lambda_{i}$ is the $i$-th singular value, $K$ is the number of samples and $\kappa$ the number of retained modes. Following Abadía-Heredia et al. (2022), the $20$ most energetic modes are retained to perform the forecasting in the laminar flow. Note in Figure 8 these number of modes contained more than $90\%$ of the overall energy.
@@
Once the models have been trained, they are tasked to iteratively predict $200$ snapshots ahead in time (from $400$ to $599$) and compare the predicted snapshots with the ground truth. This comparison is done at two representative spatial coordinates shown in Figure 9, where Figure 5 presents the predictions obtained for the streamwise velocity evolution from the POD-DL, residual autoencoder, and variational autoencoder models, respectively. Figure 7 shows the same comparison for the wall-normal velocity evolution.
Note how in the streamwise velocity (Figure 5) the predictions from the POD-DL are the closest to the ground truth, in comparison to the other two models. A similar scenario can be observed for the evolution of the wall-normal velocity (Figure 7), where predictions from the POD-DL model are also the most accurate, with respect to the reference. Nonetheless, predictions from the residual autoencoder are more accurate than the VAE model.
Figure 9: Snapshots from the laminar flow streamwise velocity, at time instants (a) $400$ and (b) $440$, representing by red points, $P_{1}$ (left) and $P_{2}$ (right), the spatial coordinates used to compare the streamwise, $u_{x}$, and wall-normal, $u_{y}$, velocity evolution shown in Figs. 5 and 7.
To evaluate the accuracy of the predictions, both the RRMSE, defined in eq. 6, and the SSIM, defined in eq. 7, are used. The SSIM measures the similarity between two figures based on their structural information. Figure 10 shows the measures obtained by these two metrics for the laminar flow. Following the results in Figures 5 and 7, the most accurate predictions, according to these metrics, are obtained from the POD-DL model. As the predictions are generated iteratively, it is intuitive to think that accumulated error will make predictions further out in time worse. However, the POD-DL model is found to be more stable over time.
This stability can also be qualitatively observed in Figs. 11 and 12, where representative snapshots are plotted comparing the ground truth and predictions from the three models. In particular, at time step $599$, predictions from the residual autoencoder and variational autoencoder show noticeable shifts, whereas predictions from the POD-DL model, although still shifted, are less so.
Figure 10: Error committed by the models, in each prediction, measured by the RRMSE in the (a) streamwise and (b) wall-normal velocity component, respectively, and the SSIM in the (c) streamwise and (d) wall-normal velocity component, respectively.
@@
(a) Ground truth snapshot and predictions at time instant $400$.
(b) Ground truth snapshot and predictions at time instant $460$.
(c) Ground truth snapshot and predictions at time instant $599$.
Figure 11: Snapshots at some representative time instants showing the ground truth streamwise velocity component and the predictions obtained from the different models. The time instants chosen are $400$, $460$ and $599$.
(a) Ground truth snapshot and predictions at time instant $400$.
(b) Ground truth snapshot and predictions at time instant $460$.
(c) Ground truth snapshot and predictions at time instant $599$.
Figure 12: Snapshots at some representative time instants showing the ground truth wall-normal velocity component and the predictions obtained from the different models. The time instants chosen are $400$, $460$ and $599$.
@@ III.2 Three-dimensional wake behind a circular cylinder: turbulent flow
This dataset is formed by a set of two-dimensional velocity fields obtained from an experimental analysis of a three-dimensional wake past a circular cylinder of diameter $D=5$. These experiments were carried out in the L10 low-speed wind tunnel of the von Karman Institute, using time-resolved particle image velocimetry (TR-PIV). The full dataset covers about $4.5$ time units, where the velocity of the free stream $U$ evolves through two steady state conditions, and the transition between them covers approximately $1$ time unit. The full dataset is formed by $13200$ snapshots, where the first $4000$ belong to the first steady state at $\hbox{Re}=4000$, the next $4000$ belongs to the transitory state and the last $5200$ belong to the second steady state at $\hbox{Re}=2600$ Mendez et al. (2020).
@@
Examining Figure 13, it is evident that the flow exhibits significantly greater complexity compared to the laminar flow shown in Figure 6. This complexity is attributable to the high Reynolds numbers at which the flow was investigated.
Through extensive trials, it was determined that the optimal approach to deal with this dataset is by simplifying the dynamics via singular value decomposition (SVD). Specifically, we reconstruct the flow by considering only the most energetic modes that encapsulate the principal dynamics. Figure 14 shows the decay of the $100$ most energetic modes and cummulative energy $E$. The cummulative energy is defined as in 10.
The optimal training performance for both the residual and variational autoencoders was achieved when retaining the $20$ most energetic modes. Table 4 illustrates the visual differences between the original flow data, derived from experimental measurements, and the flow reconstruction obtained using singular value decomposition (SVD) with these 20 dominant modes. As shown in Figure 14, this set of modes captures approximately $23\%$ of the total energy.
The simplified dataset, retaining only the 20 most energetic modes, is utilized to train both the residual and variational autoencoders. In contrast, the POD-DL model, which inherently applies SVD, is trained on the original dataset. However, for consistency with the autoencoders, the POD-DL model is also constrained to the $20$ most energetic modes. In all cases, SVD is applied exclusively to the initial $2800$ samples used for training the models.
@@
Once the models are trained, they are tasked to iteratively predict $200$ snapshots ahead in time (from snapshots [$2801$ $3000$]) and compare the statistics of the predicted snapshots with the ground truth snapshots obtained from experimental measurements. Figure 15 presents this comparison, at two representative spatial coordinates shown in Figure 16, where left, middle and right columns depict the streamwise velocity predictions from the POD-DL, residual autoencoder, and variational autoencoder (VAE) models, respectively.
@@
Figure 15: Comparison of the ground truth turbulent flow streamwise velocity evolution $u_{x}(t)$, at the spatial coordinate $P_{1}$ shown in Figure 16, against the predictions obtained from the (a) POD-DL, (b) residual autoencoder and (c) variational autoencoder models, respectively. Counter part for the spatial coordinate $P_{2}$ in figures (d), (e) and (f). In all figures the ground truth is represented by a solid line and the prediction by a dashed line.
Figure 16: Snapshots from the turbulent flow streamwise velocity, at time instants (a) $2801$ and (b) $2850$, representing by red points, $P_{1}$ (left) and $P_{2}$ (right), the spatial coordinates used to compare the streamwise, $u_{x}$ velocity evolution shown in Figure 15.
@@
The analysis confirms our earlier assumption based on the streamwise velocity comparison: the POD-DL model yields the most accurate predictions, as evidenced by both the RRMSE and SSIM metrics. This is crucial to note because while the RRMSE values in Figure 17 (a) suggest that all models produce predictions of similar accuracy, the SSIM values in Figure 17 (b) reveal a stark difference. Specifically, the SSIM shows that the predictions from the VAE are significantly less similar to the ground truth data compared to the other models. This discrepancy arises because the SSIM considers structural information, providing a more nuanced assessment of similarity between two snapshots. This highlights the importance of using complementary metrics like SSIM alongside RRMSE to obtain a comprehensive evaluation of model performance.
The pronounced differences in prediction accuracy among the models are further corroborated in Figure 19, where visual comparisons of the streamwise velocity snapshots demonstrate the superior performance of the POD-DL model.
@@
In Figure 18, we compare the velocity magnitude histograms of the real snapshots with the predictions from the three models. The results indicate that predictions from the POD-DL model exhibit a distribution that is closer to the ground truth data compared to the other two models, specially the variational autoencoder whose predictions are mostly the mean flow with noise, as shown in Figure 19. This observation aligns with the earlier findings presented in Figs. 15 and Figure 17, further validating the superior performance of the POD-DL model in capturing the statistical properties of the turbulent flow.
@@
(a) Ground truth snapshot and predictions at time instant $2801$.
(b) Ground truth snapshot and predictions at time instant $2860$.
(c) Ground truth snapshot and predictions at time instant $3000$.
Figure 19: Snapshots at some representative time instants showing the ground truth streamwise velocity component and the predictions obtained from the different models. The time instants chosen are $2801$, $2860$ and $3000$. Note for this dataset the dimensions of the snapshots are in millimeters (mm).

### v4

@@ II.2 Probabilistic forecasting: Variational autoencoder
In contrast to point forecasting, our goal here is to approximate the probability distribution $\mathbb{F}_{\bm{v}}$ that governs the forecasting process from a sequence of previous snapshots $\{\bm{v}_{i-n+1},\bm{v}_{i-n+2},\dots,\bm{v}_{i}\}$ to the next snapshot $\{\bm{v}_{i+1}\}$. To achieve this, we define a third model which adopts the Bayesian perspective of statistical inference and develop a variational inference model (Blei, Kucukelbir, and McAuliffe, 2017) based on variational autoencoders (VAEs) (Kingma and Welling, 2022). This model is presented in this article for the first time, to the authors’ knowledge. For a visual depiction of the proposed model, see Fig. 3.
Figure 3: Overview of the architecture of the VAE. Similar to the residual autoencoder, the encoder and decoder of this model are composed of convolutional neural networks.
As with the residual autoencoder, VAE adopts an encoder-decoder structure that operate on the spatial dimension of the snapshots, and there is a hidden block that operates on the temporal dimension to capture the temporal correlations within the latent space. The specific architecture used for the laminar flow is detailed in Tables 10 and 11.
@@
Note from Fig. 3 that the decoder input comes from the posterior distribution $p_{0}(\bm{z}|\{\bm{v}_{i}\}_{i=t-n+1}^{t})$. Therefore, it is necessary to generate a sample from this posterior and use it as input to the decoder. The same is done with the likelihood $p(\bm{v}_{t+1}|\bm{z})$ to obtain the time-ahead predicted snapshot.
@@
Where $\bm{x}$ is the input image, $p_{0}(\bm{z}|\bm{x})$ is the exact posterior of the latent variables $\bm{z}$, $q(\bm{z}|\bm{x})$ is the approximate posterior, $p(\bm{x}|\bm{z})$ is the likelihood of the image $\bm{x}$ given the latent variables $\bm{z}$ and KL[||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] is the Kullback-Leibler divergence. Note that in the original formulation (5), the ELBO is evaluated for a single image $\bm{x}$.
In this work, we explore a modification of this loss function (5), where the exact and approximate posterior are conditioned on the input sequence of snapshots $\{\bm{v}_{i}\}_{i=t-n+1}^{t}$ instead of a single image $\bm{x}$. Additionally, we aim to approximate the likelihood of the time-ahead snapshot $p(\bm{v}_{t+1}|\bm{z})$. Therefore the loss function used to train the VAE is as follows,
@@
This newly defined loss function allows to approximate the likelihood of the time-ahead snapshot based on the latent variables $\bm{z}$ that are obtained from the previous sequence of snapshots $\{\bm{v}_{i}\}_{i=t-n+1}^{t}$. Since this study focuses on a forecasting problem, the loss function used to train the VAE is the ELBOf (6), rather than the standard ELBO (5).
@@ III Results
After extensive experimentation, it was observed that the residual autoencoder and VAE models perform better when tasked with predicting a single time-ahead sample, in contrast to the POD-DL model. The architectural simplicity of the POD-DL model enables it to leverage a smaller training set while still predicting multiple time-ahead predictions effectively, i.e., a longer horizon.
The number of training and testing samples for each dataset is summarized in Table 3. These values represent the optimal training sample sizes determined for the residual autoencoder and VAE models. However, Appendix B presents an analysis of the POD-DL model’s performance when trained with varying numbers of samples, highlighting its behavior under different training set sizes.
The selection of hyperparameters for all models was informed by extensive experimentations, aiming to balance the model complexity with training efficiency. For the residual autoencoder, hyperparameter tuning prioritized optimizing the capture of both spatial and temporal dependencies, as a smaller latent dimension may simplify forecasting but risks insufficient representation of spatial dependencies, potentially degrading the accuracy of dynamic predictions. Effective forecasting hinges on the encoder’s ability to accurately capture the spatial structures within the flow, necessitating a careful trade-off in the latent dimension size.
For the VAE, hyperparameter tuning followed a similar approach to the residual autoencoder, with an additional focus on balancing the reconstruction loss and the Kullback-Leibler divergence.
Finally, in the case of the POD-DL model, hyperparameter tuning was centered on selecting an appropriate number of POD modes. This process aimed to capture the most energetically significant structures within the flow dynamics while ensuring computational feasibility, since the greater the number of modes retained, the longer and more complex the training will be. This is because the less energetic modes has a noisier dynamics. Truncating to the optimal number of modes was critical to preserving essential dynamics and maintaining a balance between model accuracy and efficiency.
@@
Figure 5: Comparison of the ground truth streamwise velocity evolution $u_{x}(t)$, at spatial coordinate $P_{1}$ shown in Fig. 9, against the predictions obtained from the (a) POD-DL, (b) residual autoencoder and (c) variational autoencoder models, respectively. Counter part for the spatial coordinate $P_{2}$ in figures (d), (e) and (f). In all figures the ground truth is represented by a solid line and the prediction by a dashed line.
@@ III.1 Three-dimensional wake behind a circular cylinder: laminar flow
This dataset represents a three-dimensional wake of a circular cylinder, from a viscous, incompressible, and Newtonian flow, which is described by the Navier–Stokes equations, written in non-conservative form, as follows,
@@
The first $100$ snapshots of the simulation correspond to the transitory state of the flow, as is shown in Fig. 6. It is known that in this regime the flow statistics are constantly varying, thus these first snapshots are not considered for training and testing the models. Figure 6 shows how snapshots from $100$ to $399$ are used to train the models, while the rest ($400$ to $599$) are used to testing. Note how these sets are disjoint allowing to test the models on unseen data during training, in this way the models are evaluated on their extrapolation ability to predict long-term temporal dynamics accurately.
@@
The decay of the $50$ most energetic modes and cummulative energy $E$ are shown in Fig. 8. The cummulative energy is defined as follows,
@@
where $\lambda_{i}$ is the $i$-th singular value, $K$ is the number of samples and $\kappa$ the number of retained modes. Following Abadía-Heredia et al. (2022), the $20$ most energetic modes are retained to perform the forecasting in the laminar flow. Note in Fig. 8 these number of modes contained more than $90\%$ of the overall energy.
@@
Once the models have been trained, they are tasked to iteratively predict $200$ snapshots ahead in time (from $400$ to $599$) and compare the predicted snapshots with the ground truth. This comparison is done at two representative spatial coordinates shown in Fig. 9, where Fig. 5 presents the predictions obtained for the streamwise velocity evolution from the POD-DL, residual autoencoder, and variational autoencoder models, respectively. Figure 7 shows the same comparison for the wall-normal velocity evolution.
Note how in the streamwise velocity (Fig. 5) the predictions from the POD-DL are the closest to the ground truth, in comparison to the other two models. A similar scenario can be observed for the evolution of the wall-normal velocity (Fig. 7), where predictions from the POD-DL model are also the most accurate, with respect to the reference. Nonetheless, predictions from the residual autoencoder are more accurate than the VAE model.
Figure 9: Snapshots from the laminar flow streamwise velocity, at time instants (a) $400$ and (b) $440$, representing by red points, $P_{1}$ (left) and $P_{2}$ (right), the spatial coordinates used to compare the streamwise, $u_{x}$, and wall-normal, $u_{y}$, velocity evolution shown in Fig. 5 and Fig. 7.
To evaluate the accuracy of the predictions, both the RRMSE, defined in eq. 7, and the SSIM, defined in eq. 8, are used. The SSIM measures the similarity between two figures based on their structural information. Figure 10 shows the measures obtained by these two metrics for the laminar flow. Following the results in Fig. 5 and Fig. 7, the most accurate predictions, according to these metrics, are obtained from the POD-DL model. As the predictions are generated iteratively, it is intuitive to think that accumulated error will make predictions further out in time worse. However, the POD-DL model is found to be more stable over time.
This stability can also be qualitatively observed in Fig. 11 and Fig. 12, where representative snapshots are plotted comparing the ground truth and predictions from the three models. In particular, at time step $599$, predictions from the residual autoencoder and variational autoencoder show noticeable shifts, whereas predictions from the POD-DL model, although still shifted, are less so.
Figure 10: Prediction error committed by the models, in the laminar flow, measured by the RRMSE in the (a) streamwise and (b) wall-normal velocity component, respectively, and the SSIM in the (c) streamwise and (d) wall-normal velocity component, respectively.
@@
(a) Ground truth snapshot, predictions and absolute error at time instant $400$.
(b) Ground truth snapshot, predictions and absolute error at time instant $460$.
(c) Ground truth snapshot, predictions and absolute error at time instant $599$.
Figure 11: Snapshots at some representative time instants showing the ground truth streamwise velocity component of the laminar flow, the predictions obtained from the different models and their respective absolute error in gray-scale. The time instants chosen are $400$, $460$ and $599$.
(a) Ground truth snapshot, predictions and absolute error at time instant $400$.
(b) Ground truth snapshot, predictions and absolute error at time instant $460$.
(c) Ground truth snapshot, predictions and absolute error at time instant $599$.
Figure 12: Snapshots at some representative time instants showing the ground truth wall-normal velocity component of the laminar flow, the predictions obtained from the different models and their respective absolute error in gray-scale. The time instants chosen are $400$, $460$ and $599$.
@@ III.2 Three-dimensional wake behind a circular cylinder: turbulent flow
This dataset is derived from an experimental analysis of a three-dimensional flow past a circular cylinder with a diameter of $D=5$. However, it consists of a set of two-dimensional velocity fields. These experiments were carried out in the L10 low-speed wind tunnel of the von Karman Institute, using time-resolved particle image velocimetry (TR-PIV). The full dataset covers about $4.5$ time units, where the velocity of the free stream $U$ evolves through two steady state conditions, and the transition between them covers approximately $1$ time unit. The full dataset is formed by $13200$ snapshots, where the first $4000$ belong to the first steady state at $\hbox{Re}=4000$, the next $4000$ belongs to the transitory state and the last $5200$ belong to the second steady state at $\hbox{Re}=2600$ Mendez et al. (2020).
@@
Examining Fig. 13, it is evident that the flow exhibits significantly greater complexity compared to the laminar flow shown in Fig. 6. This complexity is attributable to the high Reynolds numbers at which the flow was investigated.
Through extensive trials, it was determined that the optimal approach to deal with this dataset is by simplifying the dynamics via singular value decomposition (SVD). Specifically, we reconstruct the flow by considering only the most energetic modes that encapsulate the principal dynamics. Figure 14 shows the decay of the $100$ most energetic modes and cummulative energy $E$. The cummulative energy is defined as in 11.
The optimal training performance for both the residual and variational autoencoders was achieved when retaining the $20$ most energetic modes. Table 4 illustrates the visual differences between the original flow data, derived from experimental measurements, and the flow reconstruction obtained using singular value decomposition (SVD) with these 20 dominant modes. As shown in Fig. 14, this set of modes captures approximately $23\%$ of the total energy.
The simplified dataset, retaining only the 20 most energetic modes, is utilized to train both the residual and variational autoencoders. In contrast, the POD-DL model, which inherently applies SVD, is trained on the original dataset. However, for consistency with the autoencoders, the POD-DL model is also constrained to the $20$ most energetic modes. In all cases, SVD is applied exclusively to the initial $2800$ samples used for training the models. Similar to the laminar flow, the models are tested on unseen data during training, i.e., in this case the models are tasked to predict from snapshot $2800$.
@@
Once the models are trained, they are tasked to iteratively predict $200$ snapshots ahead in time (from snapshots [$2801$ $3000$]) and compare the statistics of the predicted snapshots with the ground truth snapshots obtained from experimental measurements. Figure 15 presents this comparison, at two representative spatial coordinates shown in Fig. 16, where left, middle and right columns depict the streamwise velocity predictions from the POD-DL, residual autoencoder, and variational autoencoder (VAE) models, respectively.
@@
Figure 15: Comparison of the ground truth turbulent flow streamwise velocity evolution $u_{x}(t)$, at the spatial coordinate $P_{1}$ shown in Fig. 16, against the predictions obtained from the (a) POD-DL, (b) residual autoencoder and (c) variational autoencoder models, respectively. Counter part for the spatial coordinate $P_{2}$ in figures (d), (e) and (f). In all figures the ground truth is represented by a solid line and the prediction by a dashed line.
Figure 16: Snapshots from the turbulent flow streamwise velocity, at time instants (a) $2801$ and (b) $2850$, representing by red points, $P_{1}$ (left) and $P_{2}$ (right), the spatial coordinates used to compare the streamwise, $u_{x}$ velocity evolution shown in Fig. 15.
@@
The analysis confirms our earlier assumption based on the streamwise velocity comparison: the POD-DL model yields the most accurate predictions, as evidenced by both the RRMSE and SSIM metrics. This is crucial to note because while the RRMSE values in Fig. 17 (a) suggest that all models produce predictions of similar accuracy, the SSIM values in Fig. 17 (b) reveal a stark difference. Specifically, the SSIM shows that the predictions from the VAE are significantly less similar to the ground truth data compared to the other models. This discrepancy arises because the SSIM considers structural information, providing a more nuanced assessment of similarity between two snapshots. This highlights the importance of using complementary metrics like SSIM alongside RRMSE to obtain a comprehensive evaluation of model performance.
The pronounced differences in prediction accuracy among the models are further corroborated in Fig. 19, where visual comparisons of the streamwise velocity snapshots demonstrate the superior performance of the POD-DL model.
@@
In Fig. 18, we compare the velocity magnitude histograms of the real snapshots with the predictions from the three models. The results indicate that predictions from the POD-DL model exhibit a distribution that is closer to the ground truth data compared to the other two models, specially the variational autoencoder whose predictions are mostly the mean flow with noise, as shown in Fig. 19. This observation aligns with the earlier findings presented in Fig. 15 and Fig. 17, further validating the superior performance of the POD-DL model in capturing the statistical properties of the turbulent flow.
@@
(a) Ground truth snapshot, predictions and absolute error at time instant $2801$.
(b) Ground truth snapshot, predictions and absolute error at time instant $2860$.
(c) Ground truth snapshot, predictions and absolute error at time instant $3000$.
Figure 19: Snapshots at some representative time instants showing the ground truth streamwise velocity component of the turbulent flow, the predictions obtained from the different models and their respective absolute error in gray-scale. The time instants chosen are $2801$, $2860$ and $3000$.